# Anisotropic Green Coordinates

### v1

@@ 6. Variational Shape Deformation of Anisotropic Green Coordinates
In Section 4, we derive the closed-form expressions for anisotropic Green coordinates. This section further demonstrates that the gradient and Hessian can also be easily obtained, forming the basis for deformation control in variational shape deformation as introduced in 2009VariationalGC and subsequent works such as 2025VariationalBiharmonic; 2024Biharmonic3D and 2025Polynomial3D. Specifically, this method defines an energy functional to determine the optimal deformation mapping that satisfies user-specified constraints while simultaneously balancing as-rigid-as-possible transformation and smoothness.
@@
For a 2D or 3D simplicial surface cage with vertices $I_{\mathbb{V}}$ and faces $I_{\mathbb{T}}$, we can define the deformation map $f_{\mathbf{a},\mathbf{b}}(\eta)$ as a linear combination of basis functions $\phi_{i}^{\mathbf{A}}(\eta)$ and $\psi_{j}^{\mathbf{A}}(\eta)$ based on the anisotropic Green coordinates as
(46) $f_{\mathbf{a},\mathbf{b}}(\eta)=\sum_{v\in I_{\mathbb{T}}}\mathbf{a}_{v}\phi_{v}^{\mathbf{A}}(\eta)+\sum_{t\in I_{\mathbb{T}}}\mathbf{b}_{t}\psi_{t}^{\mathbf{A}}(\eta),$
@@
where $\mathbf{b}_{t}=\mathbf{A}\mathbf{n}_{t}$. Here, $\mathbf{a}_{v}$ and $\mathbf{b}_{t}$ are treated as independent variables and optimized separately to expand the deformation space. Specifically, $\mathbf{a}_{v}\in\mathbb{R}^{3}$ represents the piecewise-linear coefficients defined on the cage vertices, while $\mathbf{n}_{t}\in\mathbb{R}^{3}$ denotes the piecewise-constant coefficients defined on the cage faces. Since $\mathbf{A}$ is invertible, we can directly optimize $\mathbf{b}_{t}=\mathbf{A}\mathbf{n}_{t}$, as $\mathbf{n}_{t}$ can be easily computed through $\mathbf{n}_{t}=\mathbf{A}^{-1}\mathbf{b}_{t}$.
@@
To perform variational shape deformation, we need the gradient and Hessian of $\phi_{i}^{\mathbf{A}}(\eta)$ and $\psi_{j}^{\mathbf{A}}(\eta)$. The core idea involves applying Lemma A.1 from the supplementary material to transform the deformation into an isotropic form. For the computation of the gradient and Hessian in isotropic coordinates, refer to Section B of the supplementary material.
@@
Let $C$ denote the original simplex of a cage, and let $C^{\prime}$ represent the transformed simplex obtained by left-multiplying all its vertices by $\mathbf{A}^{-1/2}$. Let $\eta_{0}$ be a point in $C$, $\mathbf{n}_{j}$ the normal of $C$, and $\mathbf{y}_{0}=\mathbf{A}^{-1/2}\eta_{0}$. We define $\psi_{j}^{\mathbf{A}}(\eta_{0},C)$ as the value of the Neumann term in anisotropic coordinates at $\eta_{0}$. Then, based on the derivations in Section 4.2, we obtain:
(47) $\psi_{j}^{\mathbf{A}}(\eta_{0},C)=\frac{1}{\sqrt{\mathbf{n}_{j}^{\top}\mathbf{A}\mathbf{n}_{j}}}\psi_{j}(\mathbf{y}_{0},C^{\prime}),$
where $\psi_{j}(\mathbf{y}_{0},C^{\prime})$ denotes the corresponding isotropic value. Then, according to Lemma A.1, the gradient $\nabla_{\eta}\psi_{j}^{\mathbf{A}}(\eta,C)$ and Hessian $H_{\eta}\psi_{j}^{\mathbf{A}}(\eta,C)$ can be calculated as:
(48) $\begin{gathered}\nabla_{\eta}\psi_{j}^{\mathbf{A}}(\eta_{0},C)=\frac{1}{\sqrt{\mathbf{n}_{j}^{\top}\mathbf{A}\mathbf{n}_{j}}}\mathbf{A}^{-1/2}\nabla_{\mathbf{y}}\psi(\mathbf{y}_{0},C^{\prime}),\\ H_{\eta}\psi_{j}^{\mathbf{A}}(\eta_{0},C)=\frac{1}{\sqrt{\mathbf{n}_{j}^{\top}\mathbf{A}\mathbf{n}_{j}}}\mathbf{A}^{-1/2}H_{\mathbf{y}}\psi(\mathbf{y}_{0},C^{\prime})\mathbf{A}^{-1/2}.\end{gathered}$
For the Dirichlet term, we denote $\phi_{i,C}^{\mathbf{A}}(\eta_{0},C)$ as the contribution of the integration over $C$ to $\phi_{i}^{\mathbf{A}}(\eta)$. Based on the derivations in Section 4.2, we have:
(49) $\phi_{i,C}^{\mathbf{A}}(\eta_{0},C)=\phi_{i,C^{\prime}}(\mathbf{y}_{0},C^{\prime}).$
@@
Then, the gradient $\nabla_{\eta}\phi_{i,C}^{\mathbf{A}}(\eta_{0},C)$ and Hessian $H_{\eta}\psi_{j}^{\mathbf{A}}(\eta,C)$ can be calculated as:
(50) $\begin{gathered}\nabla_{\eta}\phi_{i,C}^{\mathbf{A}}(\eta_{0},C)=\mathbf{A}^{-1/2}\nabla_{\mathbf{y}}\phi_{i,C^{\prime}}(\mathbf{y}_{0},C^{\prime})\\ \ H_{\eta}\phi_{i,C}^{\mathbf{A}}(\eta_{0},C)=\mathbf{A}^{-1/2}H_{\mathbf{y}}\phi_{i,C^{\prime}}(\mathbf{y}_{0},C^{\prime})\mathbf{A}^{-1/2}.\end{gathered}$
The computation of $\nabla_{\mathbf{y}}\psi(\mathbf{y}_{0},C^{\prime})$, $H_{\eta}\psi_{j}^{\mathbf{A}}(\eta,C)$$\nabla_{\mathbf{y}}\phi_{i,C^{\prime}}(\mathbf{y}_{0},C^{\prime})$ and $H_{\mathbf{y}}\phi_{i,C^{\prime}}(\mathbf{y}_{0},C^{\prime})$ can be found in Section B.2.
@@
Next, we return to the formulation of the variational deformation. We express the deformation mapping $f_{\mathbf{a},\mathbf{b}}(\eta)$ in matrix form as:
@@
where $d=2,3$. Here, $n=|I_{\mathbb{V}}|$ denotes the number of vertices, and $m=|I_{\mathbf{T}}|$ represents the number of faces. The vector $\Phi$ is a row vector composed of all Dirichlet terms $\phi_{v}^{\mathbf{A}}(\eta)$, and $\Psi$ is a row vector composed of all Neumann terms $\psi_{t}^{\mathbf{A}}(\eta)$.
@@
where each column of $\mathbf{G}_{\Phi}$ captures the gradients of $\phi_{v}^{\mathbf{A}}(\eta)$ for one vertex $v$, and each column of $\mathbf{G}_{\Psi}$ records the gradients of $\psi_{t}^{\mathbf{A}}(\eta)$ for one face $t$.
@@
where each column of $\mathbf{H}_{\Phi}$ records the elements of the Hessian matrix of $\phi_{v}^{\mathbf{A}}(\eta)$, comprising $d(d+1)/2$ values. This vector is derived by flattening the symmetric Hessian matrix, retaining the diagonal elements and the upper-right components. Similarly, $\mathbf{H}_{\Psi}$ contains the elements of the Hessian matrix of $\psi_{t}^{\mathbf{A}}(\eta)$. Since the solution to the anisotropic Laplacian equation also satisfies the boundary maximum principle, points for computing Hessians can be selected near the boundary, similar to previous methods
@@
Based on the above formulation, we can now define an optimization framework. The energy functional $E(f_{\mathbf{a},\mathbf{b}})$ is designed to be as rigid as possible (ASAP) at specific points (typically medial axis points), to ensure global smoothness (by minimizing the Hessian magnitude), and to satisfy user constraints at certain points (i.e., deforming specific points to the positions defined by the user). For discrete sampling, this problem is formulated as:
@@
(54) $\begin{gathered}\min_{\mathbf{a},\mathbf{b},\mathbf{R}_{i}}E(f_{\mathbf{a},\mathbf{b}})=\sum_{i=1}^{d}\left\|\mathbf{J}_{f}(m_{i})-\mathbf{R}_{i}\right\|_{F}^{2}+\lambda_{1}\sum_{i=1}^{r}\|f_{\mathbf{a},\mathbf{b}}(q_{i})-f_{i}\|_{2}^{2}\\ +\lambda_{2}\sum_{i=1}^{k}\left\|\mathbf{H}_{f}(w_{i})\right\|_{F}^{2}+\lambda_{3}(\|\mathbf{a}-\mathbf{a}_{0}\|_{F}^{2}+\|\mathbf{b}-\mathbf{b}_{0}\|_{F}^{2}),\\ \mathrm{s.t.},\mathbf{R}_{i}^{\top}\mathbf{R}_{i}=\mathbf{I},\forall i=1,\dots,d.\end{gathered}$
@@
Here, $\mathbf{a}_{0}$ and $\mathbf{b}_{0}$ represent the values of the source cage. The formulation here aligns more closely with 2025VariationalBiharmonic than with the earlier work (2009VariationalGC), which employs soft positional constraints $\|f_{\mathbf{a},\mathbf{b}}(q_{i})-f_{i}\|_{2}^{2}$; however, experiments also demonstrate that soft constraints effectively satisfy the requirements. We also introduce a regularization term $\|\mathbf{a}-\mathbf{a}_{0}\|_{F}^{2}+\|\mathbf{b}-\mathbf{b}_{0}\|_{F}^{2}$ with a small value of $\lambda_{3}$ to ensure minimal deviation while improving optimization stability. We apply the local/global strategy to solve this problem (2007ARAP; 2008LocalGlobal; 2009VariationalGC). Specifically, in the local step, we fix $\mathbf{a},\mathbf{b}$ and solve for the optimal rotation matrices $\mathbf{R}_{i}$. This is a well-known orthogonal Procrustes problem, which can be efficiently solved using singular value decomposition. In the global step, we fix $\mathbf{R}_{i}$ and solve for $\mathbf{a},\mathbf{b}$. This constitutes an unconstrained least-squares problem, which can be straightforwardly transformed into a linear system.
The variational shape deformation presented in this section underscores the value of establishing the anisotropic Green coordinates formulation, especially when compared to the similarity Green deformation discussed in Section 5. We denote the deformation map of isotropic Green coordinates as $f^{\prime}_{\mathbf{a}^{\prime},\mathbf{b}^{\prime}}(\eta^{\prime})$. Based on the preceding analysis, the deformation mapping can be expressed as
(55) $f_{\mathbf{a},\mathbf{b}}(\eta)=\mathbf{A}^{1/2}f^{\prime}_{\mathbf{a}^{\prime},\mathbf{b}^{\prime}}(\mathbf{A}^{-1/2}\eta).$
The gradient of $f_{\mathbf{a},\mathbf{b}}(\eta)$ can be expressed as
(56) $\nabla f_{\mathbf{a},\mathbf{b}}(\eta)=\mathbf{A}^{1/2}\nabla{f^{\prime}}(\mathbf{A}^{-1/2}\eta)\mathbf{A}^{-1/2}.$
Then, if $\mathbf{R}$ is a rotation matrix, the constraints $\nabla f_{\mathbf{a},\mathbf{b}}(\eta)=\mathbf{R}$ can be expressed as $\nabla{f^{\prime}}(\mathbf{A}^{-1/2}\eta)=\mathbf{A}^{-1/2}\mathbf{R}\mathbf{A}^{1/2}$. However, the right-hand side is not necessarily a rotation matrix, since
(57) $(\mathbf{A}^{-1/2}\mathbf{R}\mathbf{A}^{1/2})^{\top}(\mathbf{A}^{-1/2}\mathbf{R}\mathbf{A}^{1/2})=\mathbf{A}^{1/2}\mathbf{R}^{\top}\mathbf{A}^{-1}\mathbf{R}\mathbf{A}^{1/2}.$
which does not necessarily equal $\mathbf{I}$. Consequently, it is challenging to formulate the local/global problem for the similarity Green transformation.
@@ 7. Experiments
This section presents the experimental results of our proposed method. First, we focus on cage-based deformation and demonstrate the results obtained with different coefficient matrices $\mathbf{A}$. Subsequently, we showcase variational deformation outcomes under the same constraints using varying $\mathbf{A}$. All experiments are performed on a laptop equipped with an Intel CPU running at 2.40 GHz and 16GB of RAM.
We begin by describing the construction of different matrices $\mathbf{A}$ for our experiments. First, we focus on the 2D scenario. Given that every symmetric positive definite matrix $\mathbf{A}$ can be decomposed into a diagonal matrix $\Lambda$ and a rotation matrix $\mathbf{P}=\begin{pmatrix}\cos{\theta}&-\sin{\theta}\\ \sin{\theta}&\cos{\theta}\end{pmatrix}$ such that $\mathbf{A}=\mathbf{P}\Lambda\mathbf{P}^{\top}$, it suffices to specify different diagonal matrices $\Lambda$ and rotation angles $\theta$. Therefore, we can denote $\mathbf{A}=\mathcal{A}(\theta,\lambda_{1},\lambda_{2})$. Moreover, in our formulation, we may assume one of the diagonal entries of the matrix to be $1.0$, since for any $c\in\mathbb{R}^{+}$, $\nabla\cdot(\mathbf{A}\nabla u)=0$ is equivalent to $\nabla\cdot(c\mathbf{A}\nabla u)=0$. Consequently, both equations lead to the same deformation mapping. Appendix C.1 of the supplementary material provides some anisotropic matrices used in 2D deformation experiments together with the $\lambda_{1},\lambda_{2},\theta$ values.
The construction of 3D matrices follows a similar approach. The 3D diagonal matrix $\Lambda$ is defined by three parameters $\lambda_{1},\lambda_{2}$ and $\lambda_{3}$. The 3D rotation matrix $\mathbf{P}$ is parameterized by three Euler angles $\alpha,\beta,\gamma$ and can be expressed as $\mathbf{P}=\mathbf{R}_{z}(\alpha)\mathbf{R}_{y}(\beta)\mathbf{R}_{x}(\gamma)$, where

### v2

@@ 6. Variational Shape Deformation of Anisotropic Green Coordinates
In this section, we derive the expressions for the derivatives of anisotropic Green coordinates, specifically the gradients and Hessians of $\phi_{i}^{\mathbf{A}}(\eta)$ and $\psi_{j}^{\mathbf{A}}(\eta)$. These derivatives form the foundation for deformation control in the variational shape deformation framework introduced in Ben-Chen et al. (2009). In this method, an energy functional is defined to determine the optimal deformation mapping that satisfies user-specified constraints while simultaneously balancing smoothness and as-rigid-as-possible transformations. Previous work also demonstrates that the introduction of anisotropy does not conflict with our desire for rigid transformations at specific positions (Lin et al., 2024).
@@
For a simplicial surface cage with vertices $\mathbb{V}$ and faces $\mathbb{T}$, we can define the deformation map $f_{\mathbf{a},\mathbf{b}}(\eta)$ as a linear combination of basis functions $\phi_{i}^{\mathbf{A}}(\eta)$ and $\psi_{j}^{\mathbf{A}}(\eta)$ based on the anisotropic Green coordinates as
(47) $f_{\mathbf{a},\mathbf{b}}(\eta)=\sum_{v\in\mathbb{V}}\mathbf{a}_{v}\phi_{v}^{\mathbf{A}}(\eta)+\sum_{t\in\mathbb{T}}\mathbf{b}_{t}\psi_{t}^{\mathbf{A}}(\eta),$
@@
where $\mathbf{b}_{t}=\mathbf{A}\mathbf{n}_{t}$. Here, $\mathbf{a}_{v}\in\mathbb{R}^{d}$ represents the piecewise-linear coefficients defined on the cage vertices, while $\mathbf{n}_{t}\in\mathbb{R}^{d}$ denotes the piecewise-constant coefficients defined on the cage faces. To enlarge the deformation space, $\mathbf{a}_{v}$ and $\mathbf{b}_{t}$ are treated as independent variables and optimized separately. Since $\mathbf{A}$ is invertible, we can directly optimize $\mathbf{b}_{t}$, as $\mathbf{n}_{t}$ can be easily computed through $\mathbf{n}_{t}=\mathbf{A}^{-1}\mathbf{b}_{t}$.
@@
To perform variational shape deformation, it is necessary to compute the gradients and Hessians of $\phi_{i}^{\mathbf{A}}(\eta)$ and $\psi_{j}^{\mathbf{A}}(\eta)$. The core idea is to apply Lemma A.1 from the supplementary material to transform the formulation into an isotropic form. For the computation of the isotropic formulation, please refer to Appendix B.
@@
Let $t_{j}$ be a face of the cage and $t_{j}^{\prime}$ its transformation via left-multiplication of vertices by $\mathbf{A}^{-1/2}$. For any point $\eta_{0}$ on $t_{j}$ with normal $\mathbf{n}_{j}$, let $\mathbf{y}_{0}=\mathbf{A}^{-1/2}\eta_{0}$ be the corresponding point in $t_{j}^{\prime}$. Moreover, let $\psi_{t_{j}}^{\mathbf{A}}(\eta_{0})$ denote the Neumann term of $t_{j}$ for the anisotropic Green coordinates at $\eta_{0}$, while $\psi_{t_{j}^{\prime}}(\mathbf{y}_{0})$ represents the Neumann term of $t_{j}^{\prime}$ for the isotropic formulation. Based on the derivations in Section 4.2, we obtain:
(48) $\psi_{t_{j}}^{\mathbf{A}}(\eta_{0})=\frac{1}{\sqrt{\mathbf{n}_{j}^{\top}\mathbf{A}\mathbf{n}_{j}}}\psi_{t_{j}^{\prime}}(\mathbf{y}_{0}),$
@@
Therefore, according to Lemma A.1, the gradient $\nabla_{\eta}\psi_{t_{j}}^{\mathbf{A}}(\eta_{0})$ and Hessian $\mathbf{H}_{\eta}\psi_{t_{j}}^{\mathbf{A}}(\eta_{0})$ can be calculated as:
(49) $\begin{gathered}\nabla_{\eta}\psi_{t_{j}}^{\mathbf{A}}(\eta_{0})=\frac{1}{\sqrt{\mathbf{n}_{j}^{\top}\mathbf{A}\mathbf{n}_{j}}}\mathbf{A}^{-1/2}\nabla_{\mathbf{y}}\psi_{t_{j}^{\prime}}(\mathbf{y}_{0}),\\ \mathbf{H}_{\eta}\psi_{t_{j}}^{\mathbf{A}}(\eta_{0})=\frac{1}{\sqrt{\mathbf{n}_{j}^{\top}\mathbf{A}\mathbf{n}_{j}}}\mathbf{A}^{-1/2}\mathbf{H}_{\mathbf{y}}\psi_{t_{j}^{\prime}}(\mathbf{y}_{0})\mathbf{A}^{-1/2}.\end{gathered}$
For the Dirichlet term, we denote $\phi_{i,t_{j}}^{\mathbf{A}}(\eta_{0})$ as the contribution of the integration over $t_{j}$ to $\phi_{i}^{\mathbf{A}}(\eta)$. Based on the derivations in Section 4.2, we have:
(50) $\phi_{i,t_{j}}^{\mathbf{A}}(\eta_{0})=\phi_{i,t_{j}^{\prime}}(\mathbf{y}_{0}).$
Then, the gradient $\nabla_{\eta}\phi_{i,t_{j}}^{\mathbf{A}}(\eta_{0})$ and Hessian $\mathbf{H}_{\eta}\phi_{i,t_{j}}^{\mathbf{A}}(\eta)$ can be calculated as:
(51) $\begin{gathered}\nabla_{\eta}\phi_{i,t_{j}}^{\mathbf{A}}(\eta_{0})=\mathbf{A}^{-1/2}\nabla_{\mathbf{y}}\phi_{i,t_{j}^{\prime}}(\mathbf{y}_{0})\\ \ \mathbf{H}_{\eta}\phi_{i,t_{j}}^{\mathbf{A}}(\eta_{0})=\mathbf{A}^{-1/2}\mathbf{H}_{\mathbf{y}}\phi_{i,t_{j}^{\prime}}(\mathbf{y}_{0})\mathbf{A}^{-1/2}.\end{gathered}$
Thus, the computation of the derivatives of the anisotropic Green coordinates is transformed into the isotropic formulation, and can be computed by the method described in Appendix B.
@@
We now formulate the objective function for variational deformation. We express the deformation mapping $f_{\mathbf{a},\mathbf{b}}(\eta)$ in matrix form as:
@@
where $n=|\mathbb{V}|$ denotes the number of vertices, and $m=|\mathbb{T}|$ represents the number of faces. The vector $\Phi$ is a row vector composed of all Dirichlet terms $\phi_{v}^{\mathbf{A}}(\eta)$, and $\Psi$ is a row vector composed of all Neumann terms $\psi_{t}^{\mathbf{A}}(\eta)$.
@@
where each column of $\mathbf{G}_{\Phi}$ captures the gradients of $\phi_{v}^{\mathbf{A}}(\eta)$, and each column of $\mathbf{G}_{\Psi}$ records the gradients of $\psi_{t}^{\mathbf{A}}(\eta)$. Moreover,
@@
where each column of $\mathbf{H}_{\Phi}$ records the elements of the Hessian matrix of $\phi_{v}^{\mathbf{A}}(\eta)$, comprising $d(d+1)/2$ values. This vector is derived by flattening the symmetric Hessian matrix, retaining the diagonal elements and the upper-right components. Similarly, $\mathbf{H}_{\Psi}$ contains the elements of the Hessian matrix of $\psi_{t}^{\mathbf{A}}(\eta)$.
@@
Based on the above formulation, we can now define an optimization framework. The energy functional $E(f_{\mathbf{a},\mathbf{b}})$ is designed to enforce as-rigid-as-possible (ARAP) behavior at specific points (typically medial axis points), ensure global smoothness (by minimizing the Hessian magnitude), and satisfy user constraints at designated points (i.e., deforming specific points to the user-defined positions). For discrete sampling, this problem is formulated as:
@@
(55) $\begin{gathered}\min_{\mathbf{a},\mathbf{b},\mathbf{R}_{i}}E(f_{\mathbf{a},\mathbf{b}})=\sum_{i=1}^{d}\left\|\mathbf{J}_{f}(\mathbf{m}_{i})-\mathbf{R}_{i}\right\|_{F}^{2}+\lambda_{1}\sum_{i=1}^{r}\|f_{\mathbf{a},\mathbf{b}}(\mathbf{q}_{i})-\mathbf{f}_{i}\|_{2}^{2}\\ +\lambda_{2}\sum_{i=1}^{k}\left\|\mathbf{H}_{f}(\mathbf{w}_{i})\right\|_{F}^{2}+\lambda_{3}(\|\mathbf{a}-\mathbf{a}_{0}\|_{F}^{2}+\|\mathbf{b}-\mathbf{b}_{0}\|_{F}^{2}),\\ \mathrm{s.t.},\mathbf{R}_{i}^{\top}\mathbf{R}_{i}=\mathbf{I},\forall i=1,\dots,d.\end{gathered}$
@@
Here, $\mathbf{a}_{0}$ and $\mathbf{b}_{0}$ denote the values of the input source cage. This formulation aligns more closely with (Michel et al., 2025) than with the earlier work (Ben-Chen et al., 2009), which employs soft positional constraints $\|f_{\mathbf{a},\mathbf{b}}(\mathbf{q}_{i})-\mathbf{f}_{i}\|_{2}^{2}$. Experiments demonstrate that soft constraints can effectively satisfy the requirements. Since the number of optimized variables is typically much larger than the number of constraints, we incorporate a regularization term $\|\mathbf{a}-\mathbf{a}_{0}\|_{F}^{2}+\|\mathbf{b}-\mathbf{b}_{0}\|_{F}^{2}$ to enhance the optimization stability. Further details regarding parameter selection and sampling point selection will be provided in the experiment section. We apply the local/global strategy to solve Eq. (55(Sorkine and Alexa, 2007; Liu et al., 2008; Ben-Chen et al., 2009). Specifically, in the local step, we fix $\mathbf{a}$ and $\mathbf{b}$ to solve for the optimal rotation matrices $\mathbf{R}_{i}$. This step formulates a well-known orthogonal Procrustes problem, which can be efficiently solved via singular value decomposition. In the global step, we fix $\mathbf{R}_{i}$ to solve for $\mathbf{a}$ and $\mathbf{b}$, resulting in an unconstrained least-squares problem that can be directly reduced to a linear system.
@@ 7. Experiments
This section presents the experimental results of our proposed method. We first focus on cage-based deformation and demonstrate the results obtained with different coefficient matrices $\mathbf{A}$. Subsequently, we showcase variational deformation outcomes subject to the same constraints using varying $\mathbf{A}$. All experiments are on a standard laptop equipped with an Intel CPU running at 2.40 GHz and 16GB of RAM.
We begin by describing how to obtain diverse matrices $\mathbf{A}$ for our experiments, focusing initially on the 2D scenario. By the spectral theorem, any symmetric positive definite matrix $\mathbf{A}$ can be decomposed into a diagonal matrix $\Lambda=\text{diag}(\lambda_{1},\lambda_{2})$ with $\lambda_{1},\lambda_{2}>0$ and an orthogonal matrix $\mathbf{P}$ such that $\mathbf{A}=\mathbf{P}\Lambda\mathbf{P}^{\top}$. Moreover, since flipping the sign of an eigenvector leaves $\mathbf{P}\Lambda\mathbf{P}^{\top}$ unchanged, we may choose $\mathbf{P}\in SO(2)$, i.e., a rotation matrix. Therefore, we parameterize $\mathbf{P}=\begin{pmatrix}\cos{\theta}&-\sin{\theta}\\ \sin{\theta}&\cos{\theta}\end{pmatrix}$ and denote $\mathbf{A}=\mathcal{A}(\theta,\lambda_{1},\lambda_{2})$. Moreover, since for any $c\in\mathbb{R}^{+}$, the equation $\nabla\cdot(\mathbf{A}\nabla u)=0$ is equivalent to $\nabla\cdot(c\mathbf{A}\nabla u)=0$ and leads to the same deformation mapping, we can normalize one of the diagonal entries of the matrix to $1.0$ in our formulation. Appendix C.1 provides some anisotropic matrices used in 2D deformation experiments together with the $\lambda_{1},\lambda_{2},\theta$ values.
The construction of 3D matrices follows a similar approach. The 3D diagonal matrix $\Lambda$ is defined by three positive parameters $\lambda_{1},\lambda_{2}$ and $\lambda_{3}$. The 3D rotation matrix $\mathbf{P}$ is parameterized by three Euler angles $\alpha,\beta,\gamma$ and can be expressed as $\mathbf{P}=\mathbf{R}_{z}(\alpha)\mathbf{R}_{y}(\beta)\mathbf{R}_{x}(\gamma)$, where